OpenStudy (anonymous):

2+3=8, 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=?? Solve it?? HELP (;

OpenStudy (anonymous):

what is it?? from which course? are the copies right?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (dan815):

oh i fially got it xD lol

OpenStudy (dan815):

2+3=8, 2*(3+1) 3+7=27, 3*(7+2) 4+5=32, 4*(5+3) 5+8=60, 6+7=72, 7+8=??

OpenStudy (dan815):

and so on... so

OpenStudy (dan815):

7*(8+6)=56+42=98